Assuming that the government decides to address the low level of production and employment caused by COVID-19 by cutting down taxes (personal income and corporate taxes).

Explain by the use of graphs, the impact of such fiscal policy on aggregate output. In your explanation, describe the interaction between the Money market,

IS-LM and AD-AS Model.


1
Expert's answer
2020-07-15T10:03:23-0400

Effect of cutting down taxes on aggregate output

AD-AS 

The AD-AS model framework describes the aggregate demand and the short run while also including the say’s law approach which focuses on aggregate supply and the long run. The AD/AS model helps determine the gross domestic product and can be described by the graph below:

IS-LM

The investment-savings (IS) and liquidity preference-money supply (LM) is a Keynesian macroeconomic model which describes how the market for economic goods (IS) interacts with the loanable funds market (LM) or money market. The graph showcases how IS and LM curves intersect as determined by the short-run equilibrium between interest rates and output.


The graph above shows the interaction of the IS-LM curve in the goods and money market respectively. The IS curve represents equilibrium in the goods market.

Y = C(Y-T) + I(r) + G

The LM curve represents equilibrium in the money market.

M/P = L(r,Y)

The intersection determines the unique combination of Y and r that satisfies equilibrium in both markets. The impact of tax cuts as described by the IS-LM can be represented by the graph below:


The initial IS-LM equilibrium is at the point E1 (r1, Y1). A tax cuts shifts the IS curve from IS1 to IS2 establishing the equilibrium point E2 (r2, Y2). A tax cut is a fiscal policy by the government and hence affects the IS curve in the goods/commodities market. From the diagram above, consumers save (1-MPC) of the tax cut, so the initial boost in spending is smaller for ∆T than for an equal ∆G and the IS curve shifts by:

"\frac {- MPC} {1-MPC} \times ∆T"

The effects on r and Y are smaller for tax changes/cuts (∆T) than for an equal change in government expenditure (∆G)


The impact of tax cuts as described by IS-LM and AD-AS Models


Expansionary fiscal policy brought about by increase in government expenditure and/or tax cuts increases the aggregate demand.

The impact of the fiscal policy of cutting down taxes (personal income and corporate taxes) consequently helps increases the output level (Y) at each value of price as represented by the shift to the right of the aggregate demand curve (AD1 ->AD2). The overall effect will be the increase in the level of production which will additionally help create employment. With such a fiscal policy in place, the negative adversities brought about by the coronavirus pandemic will be mitigated effectively.